# Order-Aware Interactive Segmentation

### v1

@@ Abstract
Interactive segmentation aims to accurately segment target objects with minimal user interactions. However, current methods often fail to accurately separate target objects from the background, due to a limited understanding of order, the relative depth between objects in a scene. To address this issue, we propose OIS: order-aware interactive segmentation, where we explicitly encode the relative depth between objects into order maps. We introduce a novel order-aware attention, where the order maps seamlessly guide the user interactions (in the form of clicks) to attend to the image features. We further present an object-aware attention module to incorporate a strong object-level understanding to better differentiate objects with similar order. Our approach allows both dense and sparse integration of user clicks, enhancing both accuracy and efficiency as compared to prior works. Experimental results demonstrate that OIS achieves state-of-the-art performance, improving mIoU after one click by on the HQSeg44K dataset and on the DAVIS dataset as compared to the previous state-of-the-art SegNext, while also doubling inference speed compared to current leading methods. The project page is https://ukaukaaaa.github.io/OIS.
@@ 1 Introduction
Recent methods for interactive segmentation (RITM (Sofiiuk et al., 2022), SimpleClick (Liu et al., 2023), SegNext (Liu et al., 2024a), SAM (Kirillov et al., 2023), HQ-SAM (Ke et al., 2024), DynaMITe (Rana et al., 2023)) suffer from three main issues. Firstly and most importantly, current methods lack a sense of relative depth between objects in a scene, which makes it difficult for them to distinguish objects accurately. For instance, as illustrated in Figure 1, current state-of-the-art interactive segmentation methods like SegNext (Liu et al., 2024a) and HQ-SAM (Ke et al., 2024) fail to separate the target object (the gate) from the background (trees and buildings). Each object has a specific order, or relative depth from each other, corresponding to its location in 3D space (e.g., the gate is in front of the building and trees in Figure 1). However, current methods fail to incorporate this information, often leading to inaccurate segmentation masks with many false positives and false negatives (Figure 1). As a result, users are required to perform multiple interactions, i.e., redundant positive clicks to capture the entire foreground and negative clicks to exclude the background. Secondly, the encoded positive and negative clicks in current methods interact with the entire image features, causing intermingling of the foreground and background, and preventing a strong discriminative representation of the target object. This is problematic because, intuitively, negative clicks should attend only to the background regions, and positive clicks should only interact with the foreground regions, as investigated in a prior work (Cheng et al., 2024), to allow better discrimination between the target object and background. Thirdly, current methods are either computationally inefficient (methods with dense integration of prompts like RITM (Sofiiuk et al., 2022), SimpleClick (Liu et al., 2023), and SegNext (Liu et al., 2024a)) or suffer from misalignment between the image features and prompts (methods with sparse integration of prompts like SAM (Kirillov et al., 2023), HQ-SAM (Ke et al., 2024), and DynaMITe (Rana et al., 2023)).
@@ 4.4 Efficiency Analysis
Time efficiency is critical in interactive segmentation models, as users expect minimal latency to ensure smooth and responsive interactions. In Table 3, we compare our methods with leading approaches in terms of efficiency (SPC and SAT Latency) and accuracy (NoC90 and 5-mIoU). The SPC (ms) reflects the average speed per click, while SAT Latency (s) captures the overall time, accounting for both image processing and click processing.
@@ A.3 Additional related work on interactive segmentation
Early interactive segmentation methods (Boykov & Jolly, 2001; Grady, 2006; Rother et al., 2004; Gulshan et al., 2010) relied on optimization techniques to solve cost functions defined by image pixels. As deep learning became more popular, approaches began incorporating user interactions directly into neural networks (Xu et al., 2016; Sofiiuk et al., 2020; Maninis et al., 2018; Jang & Kim, 2019; Lin et al., 2020). Further advancements such as RITM (Sofiiuk et al., 2022; Liu et al., 2022) leveraged large-scale data for robustness, FocalClick (Chen et al., 2022) introduced local refinement modules, and SimpleClick (Liu et al., 2023) improved performance using vision transformers. To enhance efficiency, methods like InterFormer (Huang et al., 2023) reduced interaction time by encoding the image only once during interactions, and SegNext (Liu et al., 2024a) applied similar idea with attention mechanism to achieve high quality with low latency. Following the introduction of SAM (Kirillov et al., 2023), numerous SAM-based methods have been proposed to improve interactive segmentation in aspects such as quality (Ke et al., 2024), interaction efficiency (Zhang et al., 2023a), and high-resolution image handling (Huang et al., 2024). Variants of interactive segmentation, like multi-object segmentation (Yue et al., 2023; Rana et al., 2023), medical image segmentation (Ma et al., 2024; Wong et al., 2023), and segmentation with controllable granularity (Zhao et al., 2024), have also been proposed.
A.4 Comparison with Other 3D information Integration Method
In this section, we compare our methods to other 3D information integration methods to show that naively incorporating depth is not as effective as our order-aware attention module discussed in Sec. 3.2 and validated in Sec. 4. Following the strategy of MM-SAM (Xiao et al., 2024), which adds an extra encoder for other modalities such as lidar, depth, and thermal, we use the depth map generated by DepthAnything-V2 (Yang et al., 2024) into their additional encoder and test this configuration on the DAVIS dataset. Note that in this configuration, we disable the order-aware attention discussed in 3.2. We call this configuration DepthEncoder in Table 5, which shows the result of this setting. Our method outperforms this configuration and MM-SAM across all metrics. This indicates that formulating the 3D spatial context information into order using our specialized order-aware attention module is a better solution than directly integrating depth maps.

### v3

@@ Abstract
Interactive segmentation aims to accurately segment target objects with minimal user interactions. However, current methods often fail to accurately separate target objects from the background, due to a limited understanding of order, the relative depth between objects in a scene. To address this issue, we propose OIS: order-aware interactive segmentation, where we explicitly encode the relative depth between objects into order maps. We introduce a novel order-aware attention, where the order maps seamlessly guide the user interactions (in the form of clicks) to attend to the image features. We further present an object-aware attention module to incorporate a strong object-level understanding to better differentiate objects with similar order. Our approach allows both dense and sparse integration of user clicks, enhancing both accuracy and efficiency as compared to prior works. Experimental results demonstrate that OIS achieves state-of-the-art performance, improving mIoU after one click by on the HQSeg44K dataset and on the DAVIS dataset as compared to the previous state-of-the-art SegNext, while also doubling inference speed compared to current leading methods.
@@ 1 Introduction
Recent methods for interactive segmentation RITM (Sofiiuk et al., 2022), SimpleClick (Liu et al., 2023), SegNext (Liu et al., 2024a), SAM (Kirillov et al., 2023), HQ-SAM (Ke et al., 2024), DynaMITe (Rana et al., 2023)) suffer from three main issues. Firstly and most importantly, current methods lack a sense of relative depth between objects in a scene, which makes it difficult for them to distinguish objects accurately. For instance, as illustrated in Figure 1, current state-of-the-art interactive segmentation methods like SegNext (Liu et al., 2024a) and HQ-SAM (Ke et al., 2024) fail to separate the target object (the gate) from the background (trees and buildings). Each object has a specific order, or relative depth from each other, corresponding to its location in 3D space (e.g., the gate is in front of the building and trees in Figure 1). Current methods fail to incorporate this information, often leading to inaccurate segmentation masks with many false positives and false negatives (Figure 1). As a result, users are required to perform multiple interactions, i.e., redundant positive clicks to capture the entire foreground and negative clicks to exclude the background. Secondly, the encoded positive and negative clicks in current methods interact with the entire image features, causing intermingling of the foreground and background, and preventing a strong discriminative representation of the target object. This is problematic because, intuitively, negative clicks should attend only to the background regions, and positive clicks should only interact with the foreground regions, as investigated in a prior work (Cheng et al., 2024), to allow better discrimination between the target object and background. Thirdly, current methods are either computationally inefficient (methods with dense integration of prompts like RITM (Sofiiuk et al., 2022), SimpleClick (Liu et al., 2023), and SegNext (Liu et al., 2024a)) or suffer from misalignment between the image features and prompts (methods with sparse integration of prompts like SAM (Kirillov et al., 2023), HQ-SAM (Ke et al., 2024), and DynaMITe (Rana et al., 2023)).
@@ 4.4 Efficiency Analysis
Time efficiency is critical in interactive segmentation models, as users expect minimal latency to ensure smooth and responsive interactions. In Table 3, we compare our methods with leading approaches in terms of efficiency (SPC and SAT Latency) and accuracy (NoC90 and 5-mIoU). The SPC (ms) reflects the average speed per click, while SAT Latency (s) captures the overall time, accounting for image processing, depth model prediction, click processing, and segmentation.
@@ 5 Conclusion
Ethics Statement. We have utilized publicly available datasets for training and evaluation, ensuring compliance with all relevant data usage policies and privacy regulations. All datasets used are anonymized to protect individual privacy, and no personally identifiable information is included in our model or results. The model is designed for general-purpose image segmentation tasks; however, we recognize that its deployment in specific applications may raise ethical considerations. We advocate for the responsible use of our technology to prevent misuse in contexts that could infringe on privacy or other ethical standards. Besides, all methodologies and experiments have been conducted with integrity, adhering to established research ethics guidelines. We are committed to transparency and reproducibility, providing thorough documentation of our processes to facilitate independent verification and ethical scrutiny. Finally, we recognize the importance of ongoing dialogue regarding the ethical implications of advancements in image segmentation technologies and encourage the research community to engage in continuous evaluation to uphold ethical standards in future developments.
Reproducibility Statement. To ensure the reproducibility of our work, we extensively describe the implementation details, such as the pre-trained model, training data, and other important training hyperparameters in Sec 3, Sec 4.1 and Sec A.1. We will release all source code pending release approval.
@@ A.3 Additional related work on interactive segmentation
Early interactive segmentation methods (Boykov & Jolly, 2001; Grady, 2006; Rother et al., 2004; Gulshan et al., 2010) relied on optimization techniques to solve cost functions defined by image pixels. As deep learning became more popular, approaches began incorporating user interactions directly into neural networks (Xu et al., 2016; Sofiiuk et al., 2020; Maninis et al., 2018; Jang & Kim, 2019; Lin et al., 2020). Further advancements such as RITM (Sofiiuk et al., 2022; Liu et al., 2022) leveraged large-scale data for robustness, FocalClick (Chen et al., 2022) introduced local refinement modules, and SimpleClick (Liu et al., 2023) improved performance using vision transformers. OACE (Mathur et al., 2024) is proposed to improve the foreground distinction in a contrastive learning manner. To enhance efficiency, methods like InterFormer (Huang et al., 2023) reduced interaction time by encoding the image only once during interactions, and SegNext (Liu et al., 2024a) applied similar idea with attention mechanism to achieve high quality with low latency. Following the introduction of SAM (Kirillov et al., 2023), numerous SAM-based methods have been proposed to improve interactive segmentation in aspects such as quality (Ke et al., 2024), interaction efficiency (Zhang et al., 2023a), and high-resolution image handling (Huang et al., 2024). Variants of interactive segmentation, like multi-object segmentation (Yue et al., 2023; Rana et al., 2023), medical image segmentation (Ma et al., 2024; Wong et al., 2023), and segmentation with controllable granularity (Zhao et al., 2024), have also been proposed.
A.4 Comparison with other interactive segmentation methods incorporating 3D information
In this section, we compare our methods to other interactive segmentation methods that integrate 3D information (MM-SAM (Xiao et al., 2024)) to show that naively incorporating depth is not as effective as our order-aware attention module discussed in Sec. 3.2 and validated in Sec. 4. Here, we take the depth maps generated from DepthAnything V2 (Yang et al., 2024a) as an input to MM-SAM and test this on the DAVIS dataset; we call this configuration MM-SAM in Table 5. Moreover, following the strategy of MM-SAM, which adds an extra encoder for other modalities such as lidar, depth, and thermal, we also train an additional Depth Encoder and add to our pipeline. Note that in this configuration, we disable the order-aware attention discussed in 3.2 to ensure fair comparison. We call this configuration DepthEncoder in Table 5. Our method outperforms this configuration and MM-SAM across all metrics. This indicates that incorporating 3D information into the model through order using our specialized order-aware attention module is a better solution than directly integrating depth maps.
@@ A.8 Failure Cases
A.9 Impact of Depth Map on Model Performance
To evaluate the impact of the quality of the depth map used on our model’s performance, we adopt three commonly used depth prediction models: DepthAnything V2 (Yang et al., 2024b), DepthAnything V1 (Yang et al., 2024a), and ZoeDepth (Bhat et al., 2023). As illustrated in Fig. 15, DepthAnything V2 produces the most fine-grained details, such as the girl’s fingers and animal’s tail, while the other two methods generate lower-quality depth maps with some possible depth prediction errors. Note that these three models generate depth maps in different quality levels, allowing us to get better insights on our model’s robustness to the depth quality.
| Methods | Depth Model | NoC90 ↓ | 5-mIoU ↑ |
| SegNext (Liu et al., 2024a) | - | 4.43 | 91.87 |
| OIS | DepthAnythingV2 (Yang et al., 2024b) | 3.80 | 92.76 |
| OIS | DepthAnythingV1 (Yang et al., 2024a) | 3.78 | 92.69 |
| OIS | ZoeDepth (Bhat et al., 2023) | 3.75 | 92.75 |
We compared our model’s performance on the DAVIS (Perazzi et al., 2016) dataset using depth maps from these three models. The results, presented in Table 6, show minimal performance variation, with our method consistently outperforming the current SOTA method, SegNext (Liu et al., 2024a). This is because the depth map is used solely to generate the order map, which guides the model’s understanding of relative depth between objects. Even a lower quality depth map (DepthAnything V1) has little impact on the final segmentation performance. More importantly, our proposed object-aware attention module is able to negate the effects of erroneous order maps caused due to erroneous depth maps.
Qualitative results are provided in Fig. 15. It is observed that even though the qualities of the depth map are different, the prediction masks are nearly the same, especially the regions that depth maps have significant different (girls’ hand and animal’s tail), which indicates the robustness of our model with different depth map source.
Moreover, it is important for the model to remain robustness when meeting the depth prediction errors. In Fig. 16, we put two examples which contain depth prediction error. In the first case, the dancer’s hat is incorrectly blending into the background audience in the predicted depth map. However, our model correctly recovers the error and accurately segments the hat. Additionally, in the second case, the duck’s boundary closely mixes with the neighboring water in terms of depth, while our model successfully separates the duck from the water in the segmentation prediction. This robustness is attributed to our proposed object-aware understanding module which ensures that our model comprehends the object as a whole, enabling it to handle depth prediction errors effectively.
Besides, it is noted that our model also achieve great performance in scenarios where significant depth variation exists within the target object. As demonstrated in Fig. 17, the target objects exhibit considerable depth variation, yet our model consistently delivers accurate and high-quality segmentation. This stems from the proposed object-aware understanding module plays a key role in ensuring the model perceives the target object as a whole, enabling it to handle significant depth variation effectively.
In summary, the key role of depth in our approach is to address complex scenarios where distinguishing the target from the background is challenging, as demonstrated in Fig. 1 and Fig. 4. In these scenarios, our model significantly outperforms existing methods. For cases where depth is less critical (Fig. 17), we show that our model remains robust and unaffected by potential negative effects of depth. This demonstrates that our model can effectively solve challenging cases while maintaining strong performance in standard scenarios, highlighting its superior performance and robustness.
A.10 Ablation study about the ordering of order/object-aware understanding modules
Here, we discuss the effect of the sequence order of order-aware understanding and object-aware understanding module. We conduct the comparison experiment on DAVIS (Perazzi et al., 2016) dataset. The results in Table 7 show that having the order-aware understanding module first, followed by the object-aware understanding module slightly decreased performance than our current sequence (object-aware understanding module first, followed by the order-aware understanding module), indicating the effectiveness of our sequence choice.
| Methods | NoC90 ↓ | NoC95 ↓ | 1-mIoU ↑ | 5-mIoU ↑ |
| order-aware understanding first | 3.84 | 9.41 | 85.74 | 92.49 |
| object-aware understanding first (current) | 3.80 | 8.59 | 87.29 | 92.76 |
A.11 Ablation Study on HQSeg44K dataset
To further prove the importance of each our proposed module, here we conduct the ablation experiment on HQSeg44K (Ke et al., 2024) dataset. The results, as displayed in Table 8, are consistent with the findings from the ablation study on the DAVIS dataset, as shown in Table 4, reaffirming that each proposed module plays a crucial role in enhancing overall performance.
| Methods | NoC90 ↓ | 5-mIoU ↑ |
| Full | 3.95 | 93.78 |
| w/o order | 4.87 (+0.98) | 92.49 (-1.29) |
| w/o object | 4.23 (+0.28) | 93.28 (-0.5) |
| w/o sparse | 5.23 (+1.28) | 90.80 (-2.98) |
| w/o dense | 4.97 (+1.02) | 91.75 (-2.03) |
A.12 Ablation study for order map with positive or negative clicks alone
We include an additional ablation study on DAVIS (Perazzi et al., 2016) dataset here to analyze the design of the order map. Table 9 shows that removing either the positive click order map or the negative click order map leads to a performance drop, confirming the effectiveness of combining both. Interestingly, we observe the following:
Impact of Positive Click Order Map
When only the negative click order map is used, the 1-mIoU metric decreases more significantly. This suggests that the positive click order map is particularly beneficial during the first click, as the first click is always a positive click.
Impact of Negative Click Order Map
When only the positive click order map is used, the NoC and 5-mIoU metrics see a larger decrease. This indicates that the negative click order map becomes more important as additional clicks are made. This is because subsequent clicks mainly involve adjustments and background removal, which rely heavily on negative clicks.
| Methods | NoC90 ↓ | NoC95 ↓ | 1-mIoU ↑ | 5-mIoU ↑ |
| pos+neg | 3.80 | 8.59 | 87.29 | 92.76 |
| pos | 4.36 | 9.89 | 85.68 | 92.04 |
| neg | 4.01 | 9.24 | 84.17 | 92.37 |
A.13 Ablation study of image encoder backbone
To show the robustness of our model to different backbones, we conduct an ablation study on the HQSeg44K (Ke et al., 2024) dataset with different image encoder backbone. We replace the DepthAnything V2 backbone (Yang et al., 2024b) with an MAE pretrained ViT backbone (He et al., 2022) to be consistent with prior SOTA methods, SegNext (Liu et al., 2024a), InterFormer (Huang et al., 2023), and SimpleClick (Liu et al., 2023). Table 10 shows that our method still significantly outperforms the other methods with MAE ViT backbone. Furthermore, Table 11 highlights that the performance gains from our proposed order and object-aware attention far exceed those from switching backbones, representing the effectiveness and importance of our proposed methods.
| Methods | Backbone | NoC90 ↓ | NoC95 ↓ | 5-mIoU ↑ |
| SimpleClick (Liu et al., 2023) | MAE ViT-B | 7.47 | 12.39 | 85.11 |
| InterFormer (Huang et al., 2023) | MAE ViT-B | 7.17 | 10.77 | 82.62 |
| SegNext (Liu et al., 2024a) | MAE ViT-B | 5.32 | 9.42 | 91.75 |
| OIS | MAE ViT-B | 4.41 | 8.01 | 93.12 |
| Methods | Backbone | NoC90 ↓ | NoC95 ↓ | 5-mIoU ↑ |
| OIS w/o order+object | MAE ViT-B | 5.54 | 9.57 | 90.58 |
| OIS | MAE ViT-B | 4.41 | 8.01 | 93.12 |
| OIS w/o order+object | DepthAnythingV2 ViT-B | 5.23 | 8.91 | 90.80 |
| OIS | DepthAnythingV2 ViT-B | 3.95 | 7.50 | 93.78 |
A.14 Performance Comparison Trained on COCO
Since tradional interactive segmentation methods, including RITM (Sofiiuk et al., 2022), FocalClick (Chen et al., 2022), SimpleClick (Liu et al., 2023), and InterFormer (Huang et al., 2023) are purely trained on COCO+LVIS dataset (Lin et al., 2014; Gupta et al., 2019), to have a fair comparison with them, we also train our model using only COCO+LVIS dataset. In Table 12, we provide the comparison results on HQSeg44K (Ke et al., 2024) and DAVIS (Perazzi et al., 2016) dataset. The results demonstrate that our method still outperforms other methods by a large margin, which indicates the effectiveness of our model.
| HQSeg44K | DAVIS | |||||
| NoC90 ↓ | NoC95 ↓ | 1-mIoU ↑ | NoC90 ↓ | NoC95 ↓ | 1-mIoU ↑ | |
| RITM (Sofiiuk et al., 2022) | 10.01 | 14.58 | 36.03 | 5.34 | 11.45 | 72.53 |
| FocalClick (Chen et al., 2022) | 7.03 | 10.74 | 61.92 | 5.17 | 11.42 | 76.28 |
| SimpleClick (Liu et al., 2023) | 7.47 | 12.39 | 65.54 | 5.06 | 10.37 | 72.90 |
| InterFormer (Huang et al., 2023) | 7.17 | 10.77 | 64.40 | 5.45 | 11.88 | 64.40 |
| OIS | 5.16 | 9.18 | 85.36 | 4.41 | 9.87 | 87.21 |